The differential equation
(1-x^2)(∂^2z/∂x^2 -2xy[∂^2z/(∂x.∂y)] + (1-y^2)(∂^2z/∂y^2) + 2(∂z/∂x) +3y(∂z/∂y)=0, is elliptic for (x,y) outside the circle x^2 + y^2=1
True or false with correct explanation
1
Expert's answer
2020-10-27T18:04:04-0400

Given differential equation is

"(1-x^2)(\u2202^2z\/\u2202x^2) -2xy[\u2202^2z\/(\u2202x.\u2202y)] + (1-y^2)(\u2202^2z\/\u2202y^2) + 2(\u2202z\/\u2202x) +3y(\u2202z\/\u2202y)=0" .

Compare it with

"A(\u2202^2z\/\u2202x^2)+B[\u2202^2z\/(\u2202x.\u2202y)] + C(\u2202^2z\/\u2202y^2) + D(\u2202z\/\u2202x) +E(\u2202z\/\u2202y)=0" , we have

"A= 1-x^2, B = -2xy, C =1-y^2" .

Now, "B^2-4AC = 4x^2 y^2 - 4(1-x^2)(1-y^2) = 4(x^2+y^2-1)"

For (x,y) outside the circle "x^2 + y^2=1" , we have "x^2 + y^2 > 1" .

"\implies B^2-4AC > 0" .

Hence, the given differential equation is Hyperbolic equation.

Thus, given statement is false.